# Spectral Variance in a Stochastic Gravitational Wave Background From a Binary Population

### v2

@@ Abstract
A population of compact object binaries emitting gravitational waves that are not individually resolvable will form a stochastic gravitational wave signal. While the expected spectrum over population realizations is well known from Phinney (2001), its higher order moments have not been fully studied before or computed in the case of arbitrary binary evolution. We calculate analytic scaling relationships as a function of gravitational-wave frequency for the statistical variance, skewness, and kurtosis of a stochastic gravitational-wave signal over population realizations. If the time derivative of the binary orbital frequency can be expressed as a power-law in frequency, we find that these moment quantities also take the form of power-law relationships. We also develop a numerical population synthesis framework against which we compare our analytic results, finding excellent agreement. These new scaling relationships provide physical context to understanding spectral fluctuations in a gravitational-wave background signal and may provide additional information that can aid in explaining the origin of the nanohertz-frequency signal observed by pulsar timing array campaigns.
@@ 1 Introduction
With strong evidence for a nanohertz-frequency gravitational wave (GW) background (GWB) signal recently presented by international collaborations of pulsar timing arrays (PTAs, Agazie et al., 2023a; Antoniadis et al., 2023a; Reardon et al., 2023; Xu et al., 2023), attention is now shifting to the elusive nature of its origin. While a population of GW-radiating supermassive black-hole binaries (SMBHBs) seems the most likely, many models of early-Universe processes can be formed to explain the PTA data (e.g. Agazie et al., 2023b; Afzal et al., 2023; Antoniadis et al., 2023b; Smarra et al., 2023).
@@
However, there has been little work to compute the expected variance or higher moments of the GWB signal for a finite population of GW sources. This could be applied to model selection and parameter estimation of the GWB to improve upon the simplistic power-law model for a binary population and provide additional leverage against which other signal models can be compared. There is a brief discussion of this in Jaffe & Backer (2003) in terms of the moment generating function of $h_{c}^{2}(f)$, while Sato-Polito & Zaldarriaga (2024) [^1] Sato-Polito & Zaldarriaga (2024) use techniques similar to Ginat et al. (2020) and Ginat et al. (2024). re-analyzed the NANOGrav 15-year data set (Agazie et al., 2023a, b) with a GWB spectral model that considers deviations from a power-law due to a finite population of SMBHBs. However, explicit scaling relationships for variance, skewness, and kurtosis of the distribution of characteristic strain as a function of frequency were not provided, and the calculations assumed that binaries were only evolving through GW emission in both works. Additionally, Renzini & Golomb (2024) investigated the GWB detection capabilities of LIGO-Virgo-KAGRA (LVK) given uncertainties in the source population from single event analyses. In this letter, we compute the variance and higher-order moments of the GWB for a general finite population of GW sources that create an unresolved GWB. We provide explicit frequency-scaling relationships for arbitrary binary evolution and compare directly to many numerically synthesized populations. These results are relevant not only for SMBHBs in the PTA frequency regime but for other sources of a GWB across the GW frequency spectrum, including LVK (e.g. Abbott et al., 2021) and LISA (Babak et al., 2023; Digman & Cornish, 2022; European Space Agency, 2024). One can regard this work as a generalization of Phinney (2001) to higher moments of the GWB strain distribution.
@@ 5 Discussion
PTA spectral characterization is likely too insensitive at present to accurately measure skewness and kurtosis, but these higher-order moments could be useful in the future. Meanwhile, the future LISA mission will be sensitive to several stochastic signals, such as cosmological stochastic backgrounds from first-order phase transitions, cosmic strings, and primordial black holes, a stochastic foreground from Galactic white-dwarf binaries, and potentially backgrounds from extragalactic stellar-mass black-hole binaries and extragalactic white-dwarf binaries (Babak et al., 2023; Farmer & Phinney, 2003; European Space Agency, 2024). Incorporating information on the statistics of spectral deviations from simple power-law models could help LISA analyses distinguish astrophysical from cosmological GWB signals.

### v3

# Spectral Variance in a Stochastic Gravitational-wave Background from a Binary Population

@@ Abstract
A population of compact object binaries emitting gravitational waves that are not individually resolvable will form a stochastic gravitational wave signal. While the expected spectrum over population realizations is well known from Phinney (2001), its higher order moments have not been fully studied before or computed in the case of arbitrary binary evolution. We calculate analytic scaling relationships as a function of gravitational-wave frequency for the statistical variance, skewness, and kurtosis of a stochastic gravitational-wave signal over population realizations due to finite source effects. If the time derivative of the binary orbital frequency can be expressed as a power-law in frequency, we find that these moment quantities also take the form of power-law relationships. We also develop a numerical population synthesis framework against which we compare our analytic results, finding excellent agreement. These new scaling relationships provide physical context to understanding spectral fluctuations in a gravitational-wave background signal and may provide additional information that can aid in explaining the origin of the nanohertz-frequency signal observed by pulsar timing array campaigns.
@@ 1 Introduction
With strong evidence for a nanohertz-frequency gravitational wave (GW) background (GWB) signal recently presented by international collaborations of pulsar timing arrays (PTAs, Agazie et al., 2023a; Antoniadis et al., 2023; Reardon et al., 2023; Xu et al., 2023), attention is now shifting to the elusive nature of its origin. While a population of GW-radiating supermassive black-hole binaries (SMBHBs) seems the most likely, many models of early-Universe processes can be formed to explain the PTA data (e.g. Agazie et al., 2023b; Afzal et al., 2023; Antoniadis et al., 2024; Smarra et al., 2023).
@@
However, there has been little work to compute the expected variance or higher moments of the GWB signal for a finite population of GW sources. This could be applied to model selection and parameter estimation of the GWB to improve upon the simplistic power-law model for a binary population and provide additional leverage against which other signal models can be compared. There is a brief discussion of this in Jaffe & Backer (2003) in terms of the moment generating function of $h_{c}^{2}(f)$, while Sato-Polito & Zaldarriaga (2024) [^1] Sato-Polito & Zaldarriaga (2024) use techniques similar to Ginat et al. (2020) and Ginat et al. (2024). re-analyzed the NANOGrav 15-year data set (Agazie et al., 2023a, b) with a GWB spectral model that considers deviations from a power-law due to a finite population of SMBHBs. However, explicit scaling relationships for variance, skewness, and kurtosis of the distribution of characteristic strain as a function of frequency were not provided, and the calculations assumed that binaries were only evolving through GW emission in both works. Additionally, Renzini & Golomb (2024) investigated the GWB detection capabilities of LIGO-Virgo-KAGRA (LVK) given uncertainties in the source population from single-event analyses. In this letter, we compute the variance and higher-order moments of the GWB for a general finite population of GW sources that create an unresolved GWB. While here we consider only population finiteness effects, we note that GW signal interference will in general also contribute to the variance and higher moments of the GWB. We provide explicit frequency-scaling relationships for arbitrary binary evolution and compare directly to many numerically synthesized populations. These results are relevant not only for SMBHBs in the PTA frequency regime but for other sources of a GWB across the GW frequency spectrum, including LVK (e.g. Abbott et al., 2021) and LISA (Babak et al., 2023; Digman & Cornish, 2022; Colpi et al., 2024). One can regard this work as a generalization of Phinney (2001) to higher moments of the GWB strain distribution.
@@ 5 Discussion
PTA spectral characterization is likely too insensitive at present to accurately measure skewness and kurtosis, but these higher-order moments could be useful in the future. Meanwhile, the future LISA mission will be sensitive to several stochastic signals, such as cosmological stochastic backgrounds from first-order phase transitions, cosmic strings, and primordial black holes, a stochastic foreground from Galactic white-dwarf binaries, and potentially backgrounds from extragalactic stellar-mass black-hole binaries and extragalactic white-dwarf binaries (Babak et al., 2023; Farmer & Phinney, 2003; Colpi et al., 2024). Incorporating information on the statistics of spectral deviations from simple power-law models could help LISA analyses distinguish astrophysical from cosmological GWB signals.
Finally, beyond the population finiteness effects studied here, the additional influence of GW signal interference between sources will contribute as a source of (cosmic) variance (Allen & Valtolina, 2024; Allen, 2023) and potentially to higher statistical moments of the GWB spectrum. Frequency-scaling relationships for signal interference effects, and the relative importance of population finiteness and signal interference, will be explored in future work.